An artist wants her painting to be in the shape of a golden rectangle. if the length of the painting is 36 inches, then what should be the width?
Answer by edjones(7311)

proportions of the golden rectangle=(1+sqrt(5))/2)=l/w
x/36=2/(1+sqrt(5))
x=72/(1+sqrt(5))
=22.249..."
.
Ed